A sphere and a cube of same material and same total surface
area are placed in the same evacuated space turn by turn, after
they are heated to the same temperature. Let initial rate of
cooling of the sphere and cube are R1 and R2 respectively, then
compare its initial rate of cooling:

Grade:12

2 Answers

Satish
6 years ago
Tsa of sphere = 4πr^2 and that of cube =6s^2 Given both are equal so 3s^2=2πr^2=>s/r={2π/3}^1/2Rate of cooling directly proportional to A/M
Satish
6 years ago
Tsa of sphere = 4πr^2 and that of cube =6s^2 Given both are equal so 3s^2=2πr^2=>s/r={2π/3}^1/2 Rate of cooling directly proportional to Area/mass=>R1/R2=3s/6r=1/2{s/r} =1/2(2π/3)^1/2=(π/6)^1/2